The width of a rectangle is 4 more than the length,l. write an equation to describe the perimeter in term of w.
Found 2 solutions by Maths68, mamiya:
Answer by Maths68(1474)

The width of a rectangle is 4 more than the length,l. write an equation to describe the perimeter in term of w.

Perimeter = p
Lenght = l
Width = w
Given
l+4=w
l=w-4
Perimeter = p = 2 length + 2 width
p=2l+2w
p=2(w-4)+2w
p=2w-8+2w
p=4w-8
Perimeter in term of w

Answer by mamiya(56)

The width of a rectangle is 4 more than the length,l. write an equation to describe the perimeter in term of w.
let p be the perimeter. we know that perimeter of a rectangle is p= 2(l+w)
the width is 4 more than the length, means w=4+l so, l=w-4
so p=2[(w-4)+w]
=2(2w-4)
= 4w -8

the answer is 4w-8